OpenStudy (anonymous):

Find an equation of the line containing the given points (-5,8) and (6,-4)

OpenStudy (anonymous):

y = mx+b. substitute the two points and find b and m

OpenStudy (anonymous):

solve a system 8=a*(-5)+b -4=a*6+b